# Femtoscopic study of the $\Omega\alpha$ interaction in heavy-ion collisions

### v3

@@ IV Numerical Results
By employing the scattering length and the effective range of the two models of potentials given in Table 1, $\Omega\alpha$ correlation functions is calculated by using LL formula (Eq. 7), and the results are compared with the ones from KP formula in Fig. 5 for three different source sizes $R=1,3$ and $5$ fm. Figure  5 demonstrates that for $R=1$ fm, the LL formulation produces significant different results compared with the KP formula at low momentum region. In principle, the LL formula can not be a good approximation where the source size is smaller than the interaction range (for interactions that include nuclei, it can be about $\gtrsim 3$ fm) [22]. On the other hand, the LL approximation is consistent with KP formula for relatively large source sizes, i.e. $R\geqq 3$ fm. Note that in Fig. 5, only the strong interaction is switched on.
@@
Figure 5: $\Omega\alpha$ correlation functions for three different source sizes. The solid and dash-dotted red lines are the results of the HAL QCD $\left(t/a=11\right)$ potentials with the KP (Eq. (6)) and LL (Eq. (7)) relations, respectively. The dotted and dashed black lines are the results of meson exchanges potentials with the KP and LL relations, respectively. The Coulomb interaction is switched off.
@@ V Summary and conclusions
In the present paper, the $\Omega\alpha$ potential is made from two available $\Omega N$ interactions, i.e., first principles HAL QCD HAL and developed meson exchange potentials, in the latter potential, the coupled channels effect is considered. $\Omega+\alpha$ potentials were obtained by using SFP model and they were fitted by Woods-Saxon type functions. The numerical results showed that $\Omega+\alpha$ potentials based on HAL potential is much deeper than the one based on ME potential by binding energies around $24$ and $7.5$ MeV, respectively. In both cases, the interaction ranges are about $3$ fm.
@@
I applied femtoscopy technique to predict $\Omega\alpha$ momentum correlation functions in the high-energy collisions looking for an additional and alternative source of knowledge relevant to the $\Omega N$ interaction. Employing two $\Omega\alpha$ potentials, correlation functions are calculated using KP formula for three different source sizes, $R=1,3$ fm and $5$ fm. The difference of potentials appears in the correlation functions by small source size around $1-3$ fm, with and without considering the Coulomb interactions, while for source size $R\gtrsim 5$ fm the correlation functions tended to become the same for both $\Omega\alpha$ potentials with Coulomb interactions. In all cases without Coulomb potential, the differences still remain significant at short distance. In conclusion, since the correlation functions are sensitive to $\Omega\alpha$ potentials behavior and Coulomb interactions, we could get important information about the effects of $\Omega$ particle in dense nuclear medium.
@@
Furthermore, the binding energy, scattering length and effective range were calculated by solving the Schrödinger equation using the fitted $\Omega\alpha$ potentials as the input for HAL and ME potential. Next, correlation functions were examined within the Lednicky-Lyuboshits approach and compared with the results of KP formula when Coulomb interaction is switched off. It was seen, as expected, that the LL formula by small source ($1$ fm) significantly differs from KP formula in the low-momentum region.
@@
Last but not least, in this theoretical study, the selection of source sizes $R=1,3$, and $5$ fm is based on previous studies of the two-hadron correlation function in $pp$ collisions and heavy ion collisions [22, 23]. The Koonin-Pratt formula, Eq. (6), is valid while the two correlated particles can be considered as well separated point-like particle. In the case of composite particle like $\alpha$, since there is a possibility of simultaneous formation of alpha particle the effective source size must be larger than those with single hadron emissions [38, 39, 40]. Therefore, basically we are facing a 5-body problem of two protons, two neutrons and $\Omega$ and a emergence of alpha particle and production of $\Omega{\text{-}}\alpha$ correlation take place at same time, this effect will be considered in future works. I hope that these theoretical studies could help to design experiments at FAIR [41], NICA, and J-PARC HI [42] in future.
@@ Acknowledgement
I thank Yuki Kamiya and Asanosuke Jinno for useful discussions, comments and sharing some data with me. I am grateful to the authors and maintainers of ”Correlation Analysis tool using the Schrödinger Equation” (CATS) [35], a modified version of which is used for calculations in this work. Discussions during the long-term workshop, HHIQCD2024, at Yukawa Institute for Theoretical Physics (YITP-T-24-02), were useful as I finished this work.

### v4

@@ IV Numerical Results
By employing the scattering length and the effective range of the two potential models given in Table 1, the $\Omega\alpha$ correlation function is calculated by using the LL formula (Eq. (7)), and the results are compared with the ones from the KP formula in Fig. 5 for three different source sizes of $R=1,3$ and $5$ fm. Figure  5 demonstrates that for $R=1$ fm, the LL formulation produces significant different results compared with the KP formula at low momentum region. In principle, the LL formula cannot be a good approximation where the source size is smaller than the interaction range (for interactions that include nuclei, it can be $\gtrsim 3$ fm) [22]. On the other hand, the LL approximation is consistent with the KP formula for relatively large source sizes, i.e. $R\geqq 3$ fm. Note that in Fig. 5, only the strong interaction is switched on.
@@
Figure 5: $\Omega\alpha$ correlation functions for three different source sizes. The solid and dash-dotted red lines are the results of the HAL QCD $\left(t/a=11\right)$ potentials with the KP (Eq. (6)) and the LL (Eq. (7)) relations, respectively. The dotted and dashed black lines are the results from the meson-exchange potential with the KP and the LL formulae, respectively. The Coulomb interaction is switched off in all cases.
@@ V Summary and conclusions
In the present paper, the $\Omega\alpha$ potential is obtained from two available $\Omega N$ interactions, i.e., first principles HAL QCD calculations and the meson-exchange model. For the latter potential, the coupled-channel effect is considered. $\Omega+\alpha$ potentials were obtained by using the SFP model and then fitted by a Woods-Saxon type function. The numerical results showed that $\Omega+\alpha$ potentials based on the HAL potential is much deeper than the one based on the ME potential with binding energies around $24$ and $7.5$ MeV, respectively. In both cases, the interaction ranges are about $3$ fm.
@@
I applied femtoscopy technique to predict $\Omega\alpha$ momentum correlation functions in high-energy nuclear collisions to look for an additional and alternative source of knowledge relevant to the $\Omega N$ interaction. Employing two $\Omega\alpha$ potentials, correlation functions are calculated using the KP formula for three different source sizes, $R=1,3$ fm and $5$ fm. The effect due to different potentials appear in the correlation functions for small source size around $1-3$ fm, with and without considering the Coulomb interaction, while for source size $R\gtrsim 5$ fm the correlation functions tended to become similar for both $\Omega\alpha$ potentials with the Coulomb interaction. In all cases without Coulomb potential, the differences still remain significant at short distance. In conclusion, since the correlation functions are sensitive to the behaviour of $\Omega\alpha$ potential and Coulomb interaction, we could obtain important information about the interactions of $\Omega$-particle in dense nuclear medium.
@@
Also, the binding energy, scattering length and effective range were calculated by solving the Schrödinger equation using the fitted $\Omega\alpha$ potentials with input from the HAL and the ME potentials. Furthermore, correlation functions were examined within the Lednicky-Lyuboshits approach and compared with the results of using the KP formula when Coulomb interaction is switched off. It was seen, as expected, that results in the low-momentum region from the LL formula for small source size ($1$ fm) significantly differ from those from the KP formula.
@@
Finally, in this theoretical study, the selection of source sizes $R=1,3$, and $5$ fm is based on previous studies of the two-hadron correlation function in $pp$ collisions and heavy ion collisions [22, 23]. The Koonin-Pratt formula, Eq. (6), is valid while the two correlated particles can be considered as well separated point-like particle. In the case of composite particles like $\alpha$, since there is a possibility of simultaneous formation of alpha particle the effective source size must be larger than those for the emission of any single hadron [39, 40, 41]. Therefore, we are basically facing a 5-body problem of two protons, two neutrons and $\Omega$, and the emergence of alpha particle and its correlation with the $\Omega$ take place at same time. This effect will be considered in future works. I hope that these theoretical studies could help to design experiments at FAIR [42], NICA, and J-PARC HI [43] in future.
@@ Acknowledgement
I thank Yuki Kamiya and Asanosuke Jinno for useful discussions, comments and sharing some data with me. I am grateful to the authors and maintainers of ”Correlation Analysis tool using the Schrödinger Equation” (CATS) [36], a modified version of which is used for calculations in this work. Discussions during the long-term workshop, HHIQCD2024 at Yukawa Institute for Theoretical Physics (YITP-T-24-02), were useful as I finished this work.